The price of Kenwood blender increased from ȼ150 to ȼ250. Following the increase in price, quantity demanded of the blender dropped from 500 to 400. Compute the point and arc elasticity of demand


1
Expert's answer
2021-05-06T15:59:17-0400


Given,

Initial price (P1)= ȼ150

New price (P2)= ȼ250

Initial quantity demanded (Qd1)= 500

New quantity demanded (Qd2)= 400

The Arc elasticity of demand can be calculated as:


Point elasticity of demand

"=\frac{\Delta Q}{\Delta P}\times \frac{initial\space P}{initial\space Q}"


"=\frac{(400-500)}{(250-150)}\times\frac{150}{500}"


"=\frac{-100}{100}\times\frac{150}{500}"


"=-0.3"


"|ed|=0.3"


point elasticity of demand is 0.3



"Arc\space Ed=\frac{\Delta Q}{\Delta P}\times \frac{P_1+P_2}{Q_1+Q_2}"


"\frac{(400-500)}{(250-150)}\times\frac{150+250}{400+500}"


"=\frac{-100}{100}\times \frac{400}{900}"


"=-0.44"


"|ed|=0.44"


Arc elasticity of demand is 0.44